OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (anonymous):

y = x +7 y = 2x +3 solve x+7 = 2x+3 4 = x y = 4+7 =11 check y = 2*4 +3 =11 checks answer 4,11 is the only point where the two lines intersect y = x+7 y = x+3 solve x+7 = x+3 7= 3 no solution the two lines never intersect this is the case when two equations are the same x=y 2x = 2y any point on the line is a solution www.yahooanswers.com

OpenStudy (anonymous):

Thanks